OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y − 8 = x2 − 6x?

hartnn (hartnn):

to find x-intercept,put y=0

OpenStudy (anonymous):

\[y − 8 = x^2 − 6x\] when y=0 \[0 = x^2 − 6x+8\] quadratic x=2 and x=4

OpenStudy (radar):

Rearrange the equation to read y=x^2 -6x + 8. You will note that it factorable (x-4)(x-2)=0 thus the zeroes occur at x=4 and x=2.